OpenStudy (anonymous):

Find the length of the curve. x=1+3t^2 y=4+2t^3 0

OpenStudy (zarkon):

you should have a formula...it works out nicely

OpenStudy (anonymous):

I got sqrt of 76t^2

OpenStudy (anonymous):

Oh, I see what I did wrong

OpenStudy (anonymous):

the length of the curve given parametric equations:\[L = \int\limits_{a}^{b} \sqrt{\left(\frac{dx}{dt}\right)^{2} + \left(\frac{dy}{dt}\right)^{2}}dt\]\[a \le t \le b\]

OpenStudy (anonymous):

No, I don't get it

OpenStudy (anonymous):

I got 7 as an answer, but that's not right.

OpenStudy (zarkon):

what do you get for your derivatives ?

OpenStudy (anonymous):

Not 7, 5

OpenStudy (anonymous):

dx/dt=6t dy/dt=6t^2

OpenStudy (zarkon):

good

OpenStudy (zarkon):

then you squared them and added them

OpenStudy (anonymous):

Yes

OpenStudy (zarkon):

what did you get?

OpenStudy (anonymous):

\[36t^2+36t^4\]

OpenStudy (zarkon):

all under a square root correct

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

simplify it

OpenStudy (anonymous):

6t+6t^2

OpenStudy (zarkon):

\[\sqrt{36t^2+36t^4}\] \[\sqrt{36t^2(1+t^2)}=\cdots\]

OpenStudy (anonymous):

oh, I see. \[6t \sqrt{1+t^2}\] u=1+t^2 du=2tdt

OpenStudy (zarkon):

yes...you have it

OpenStudy (anonymous):

I think I'm doing something wrong I get the answer to be 3 and the book says it's 4sqrt(2)-1

OpenStudy (zarkon):

that is what your book gives you?

OpenStudy (zarkon):

I get \[4\sqrt{2}-2\]

OpenStudy (anonymous):

It's right I checked on a graphing calc.

OpenStudy (anonymous):

Something is wrong.

OpenStudy (anonymous):

You are right

OpenStudy (anonymous):

I wrote it wrong, my bad.

OpenStudy (zarkon):

\[\int\limits_{0}^{1}6t\sqrt{1+t^2}dt\] \[u=1+t^2\] \[du=2tdt\] \[\frac{du}{2}=tdt\] \[\int\limits_{1}^{2}3\sqrt{u}du\] correct?

OpenStudy (anonymous):

Yeah

OpenStudy (zarkon):

\[=3\frac{2}{3}u^{3/2}\Big|_1^2\] \[=2u^{3/2}\Big|_1^2=2(2^{3/2})-2(1^{3/2})\]

OpenStudy (zarkon):

\[=4\sqrt{2}-2\]

OpenStudy (anonymous):

I must have put something else in because I just got 4, thanks!